### v1

@@ 4.2 Conjugating by the right matrix
$\Delta=\begin{pmatrix}\lambda_{1}&&&\\ &\lambda_{2}&&\\ &&\ddots&\\ &&&\lambda_{m}\end{pmatrix}$
@@
$f(\Delta)=\begin{pmatrix}f(\lambda_{1})&&&\\ &f(\lambda_{2})&&\\ &&\ddots&\\ &&&f(\lambda_{m})\end{pmatrix}.$
@@
  • Find the coefficients of $f(\Delta)$.
@@
We now pick a matrix $C=(c_{ij})_{i,j}\in\mathcal{C}(\Lambda)$, and a matrix $D=(d_{ij})_{i,j}\in\mathcal{D}(\Lambda)$. After multiplying by an element of $\mathbb{F}_{q}$, we may suppose that they have the same characteristic polynomial, and solve $D=f(\Delta)Cf(\Delta)^{-1}$. This means solving the system of $m(m-1)/2-(k-1)$ equations
$d_{ij}=f(\lambda_{i})f(\lambda_{j})^{-1}c_{ij}$
which yields the $m$ values $f(\lambda_{1}),\dots,f(\lambda_{m})$ up to a scalar multiple. We then find a polynomial $f$ corresponding to these values using Lagrange interpolation.
@@
Diagonal $\Delta\in\mathbb{F}_{q}^{m\times m}$ such that $\exists f\in\mathbb{F}_{q}[t]\colon\mathcal{D}=f(\Delta)\mathcal{C}f(\Delta)^{-1}$
Result: Polynomial $f\in\mathbb{F}_{q}[t]$ such that $\mathcal{D}=f(\Delta)\mathcal{C}f(\Delta)^{-1}$
@@
             Pick $C\in\mathcal{C}(\Lambda),D\in\mathcal{D}(\Lambda)$ with the same characteristic polynomial
             Solve system $d_{ij}=u_{i}u_{j}^{-1}c_{ij}$ for $(i,j)\in\{1\dots m\}^{2}$
             Compute polynomial $f$ such that $f(\delta_{ii})=u_{i}$
@@
4.3 Complexity analysis
@@
Given conjugate codes $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times m}$ with one-dimensional hulls and generators of these hulls with separable characteristic polynomials, the average complexity of finding $P\in\mathbb{F}_{q}^{m\times m}$ such that $\mathcal{D}=P\mathcal{C}P^{-1}$ using Algorithms 4 and 5 is
$\tilde{\mathcal{O}}(km^{2\omega}q^{3\sqrt{m}/(\sqrt{\log m}\log q)}),$
which is negligible compared to $q^{m/2}$.
@@
$d=\exp\left(C\sqrt{m/\log(m)}+O\left(\sqrt{m}\log(\log m)/\log(m)\right)\right)$
where $C<3$. This shows that $d=\mathcal{O}(q^{3\sqrt{m}/(\sqrt{\log m}\log q)})\ll q^{m/2}$. We can do all the computations over $\mathbb{F}_{q^{d}}$, which means the number of $\mathbb{F}_{q}$-operations will be that of $\mathbb{F}_{q^{d}}$-operations multiplied by $\tilde{\mathcal{O}}(d)$ (using FFT-based algorithm for polynomial arithmetic, see for instance [13, Thm. 8.23]). Diagonalizing $U$ and $V$ is done in time $\mathcal{O}(m^{\omega})$. Computing the subspaces $\mathcal{C}\cap E_{\Lambda}$, $\mathcal{D}\cap E_{\Lambda}$ is just linear algebra and requires $\mathcal{O}(km^{2\omega})$ operations in $\mathbb{F}_{q}$. Solving the system of equations takes $\mathcal{O}(m)$ multiplications in $\mathbb{F}_{q^{d}}$. In total, the complexity is $\tilde{\mathcal{O}}(dkm^{2\omega})$. ∎
@@ Appendix 0.A Normalizing matrices and characteristic polynomials
$\{U\in\mathbb{F}_{q}^{m\times m}\mid\operatorname{Tr}(U)=\operatorname{Tr}(U^{2})=0\}/\mathbb{F}_{q}^{\times}$ $\longrightarrow(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-1}\times\{0\}))/\mathbb{F}_{q}^{\times}$
@@
where the characteristic polynomial of $U$ is $t^{m}+a_{m-3}t^{m-3}+\dots+a_{1}+a_{0}$ and $\mathbb{F}_{q}^{\times}$ acts on $\mathbb{F}_{q}^{m-2}$ via $\lambda\diamond(a_{m-3},\dots,a_{0})=(\lambda^{3}a_{m-3},\dots,\lambda^{m}a_{0})$.
@@
The set $(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-1}\times\{0\}))/\mathbb{F}_{q}^{\times}$ has
@@
The subset $\mathrm{Sep}_{q,m}\subset(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-1}\times\{0\}))/\mathbb{F}_{q}^{\times}$ of classes of separable polynomials has $\sim q^{m-3}$ elements.
@@
elements by the Serre bound [28, Théorème]. Since every element of $\mathbb{P}^{m-3}_{3,\dots,m}(\mathbb{F}_{q})$ has exactly $q-1$ preimages in $\mathbb{F}_{q}^{m-2}$ [23, Lem. 7], this means that there are $\mathcal{O}(mq^{m-4})$ classes of inseparable polynomials in $\mathbb{P}^{m-3}_{3,\dots,m}(\mathbb{F}_{q})$. Hence, by Lemma 13, $\mathrm{Sep}_{q,m}$ has $q^{m-3}-\mathcal{O}(mq^{m-4})\sim q^{m-3}$ elements.∎
@@
      $\mathbb{F}_{q}^{\times}/\mu_{d_{0}}(\mathbb{F}_{q})^{i_{1}}=\mathbb{F}_{q}^{\times}/\mu_{d_{1}/d_{0}}(\mathbb{F}_{q})$
@@
      If $d_{1}\neq 1$, continue with $d_{2}=\gcd(d_{1},i_{2})$. Stop after the $k$-th step if $k=\ell$ or $d_{k}=1$. Return $\chi^{\prime}=\lambda_{k}\cdot\chi$. The algorithm is summed up below.
@@
while $d\neq 1\wedge(a_{m-i-1},\dots,a_{1})\neq(0,\dots,0)$ do
@@
return $\lambda$
@@
The complexity of Algorithm 6 is $\mathcal{O}(m\log(q))$. To reduce its impact on the complexity of our attack, one can precompute the normalization of the most frequent characteristic polynomials. For instance, on may compute in advance the suitable $\lambda$ for vectors $(a_{m-3},\dots,a_{0})$ such that $a_{m-3},a_{m-4}\neq 0$. Since 3 and 4 are coprime, a unique $\lambda$ is found knowing only $a_{m-3},a_{m-4}$: constructing a dictionary $\{(a_{m-3},a_{m-4}):\lambda\}$ allows to precompute the normalization for $(q-1)^{2}q^{m-4}\sim q^{m-2}$ elements of $\mathbb{F}_{q}^{m-2}$, that is, almost all of them.
@@ Appendix 0.B Proportion of codes with one-dimensional hull
Sendrier’s work [25] gives detailed results about the number of codes with a hull of given dimension. His results are proven in the case of codes inside $\mathbb{F}_{q}^{n}$ with the usual inner product. In our case however, we consider $\mathbb{F}_{q}^{m\times m}$ endowed with the bilinear form $(X,Y)\mapsto\operatorname{Tr}(XY)$. Denoting by $\sigma_{n,i}$ the number of totally isotropic $[n,k]_{q}$-codes for a given bilinear form, the number $A_{n,k,1}$-codes whose intersection with their orthogonal complement has dimension 1 is equal to [25, Theorem 2]
@@
where [${}{n}{k}$ ] is the Gaussian binomial coefficient which denotes the number of $k$-dimensional linear subspaces of $\mathbb{F}_{q}^{n}$. The proof of this result does not involve the nature of the considered non-degenerate bilinear form. Sendrier goes on to show [25, Theorem 3], using asymptotic results based on explicit values of $\sigma_{n,i}$ specific to a bilinear form of discriminant 1, that for $21\leqslant k\leqslant n/$ 2,
@@
There are different formulas of $\sigma_{n,i}$ given in [25, Theorem 1] depending on the remainders of $n,q$ modulo 2 and 4 and on the size of $k$. However, they are asymptotically equivalent, which yields this uniform result. For a bilinear form of different discriminant, these formulas are simply permuted; a general expression may be found in [8, IV, Proposition 3.5]. This does not change the asymptotic result above. Moreover, for $k\geqslant n/2$, the number of $[n,k]$-codes with one-dimensional hull is that of $[n,n-k]$-codes with one-dimensional hull, since the hull of a code $\mathcal{C}$ is exactly that of its dual. In particular, this means that for any $k$ such that $11\leqslant k\leqslant n-$ 1,

### v2

@@ 4.2.2 Conjugating by the right matrix
$\Delta=\begin{pmatrix}\delta_{1}&&&\\ &\delta_{2}&&\\ &&\ddots&\\ &&&\delta_{m}\end{pmatrix}$
@@
$f(\Delta)=\begin{pmatrix}f(\delta_{1})&&&\\ &f(\delta_{2})&&\\ &&\ddots&\\ &&&f(\delta_{m})\end{pmatrix}.$
@@
  • Find the entries of $f(\Delta)$.
@@
We now pick a matrix $C=(c_{ij})_{i,j}\in\mathcal{C}(\Lambda)$, and a matrix $D=(d_{ij})_{i,j}\in\mathcal{D}(\Lambda)$. After multiplying by an element of $\mathbb{F}_{q}$, we may suppose that they have the same characteristic polynomial, and solve $Df(\Delta)=f(\Delta)C$. This means solving the system of $m(m-1)/2-(k-1)$ equations
$d_{ij}f(\lambda_{j})=f(\lambda_{i})c_{ij}$
which yields the $m$ values $f(\delta_{1}),\dots,f(\delta_{m})$ up to a scalar multiple. We then find a polynomial $f$ corresponding to these values using Lagrange interpolation.
@@
Diagonal $\Delta=\text{Diag}(\delta_{1},\dots,\delta_{n})\in\mathbb{F}_{q}^{m\times m}$ s.t. $\delta\in\mathcal{C}$ and $\Delta\in\mathcal{D}$
Result: Polynomial $f\in\mathbb{F}_{q}[t]$ such that $\mathcal{D}=f(\Delta)\mathcal{C}f(\Delta)^{-1}$ (if exists)
@@
             Pick $C\in\mathcal{C}(\Lambda),D\in\mathcal{D}(\Lambda)$ with the same characteristic polynomial (if exist)
             Solve system $d_{ij}u_{j}=u_{i}c_{ij}$ for $(i,j)\in\{1\dots m\}^{2}$
             if The system has no solution then
                  return $\bot$
            Compute polynomial $f$ such that $f(\delta_{i})=u_{i}$
@@
Data: Codes $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times m}$
Matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$
Result: A matrix $P$ (if exists), such that $\mathcal{D}=P\mathcal{C}P^{-1}$
Compute $(\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta,S)$ using Algorithm 4 with inputs ($\mathcal{C},\mathcal{D},U,V,R$)
Compute $f$ using Algorithm 5 with inputs ($\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta$)
if $f=\bot$ then
       return $\bot$
return $Sf(\Delta)S^{-1}R$
Algorithm 6 Find the right $P$
4.2.3 Complexity analysis
@@
Given conjugate codes $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times m}$ with one-dimensional hulls and generators of these hulls with separable characteristic polynomials, the average complexity of finding $P\in\mathbb{F}_{q}^{m\times m}$ such that $\mathcal{D}=P\mathcal{C}P^{-1}$ using Algorithm 6 is
$\tilde{\mathcal{O}}(k^{\omega-1}m^{2}q^{3\sqrt{m}/(\sqrt{\log m}\log q)}).$
@@
$d=\exp\left(C\sqrt{m/\log(m)}+\mathcal{O}\left(\sqrt{m}\log(\log m)/\log(m)\right)\right)$
where $C<3$. This shows that $d=\mathcal{O}(q^{3\sqrt{m}/(\sqrt{\log m}\log q)})$. We can do all the computations over $\mathbb{F}_{q^{d}}$, which means the number of $\mathbb{F}_{q}$-operations will be that of $\mathbb{F}_{q^{d}}$-operations multiplied by $\tilde{\mathcal{O}}(d)$ (using FFT-based algorithm for polynomial arithmetic, see for instance [13, Thm. 8.23]). Diagonalizing $U$ and $V$ is done in time $\mathcal{O}(m^{\omega})$. Computing the subspaces $\mathcal{C}\cap E_{\Lambda}$, $\mathcal{D}\cap E_{\Lambda}$ is just linear algebra and requires $\mathcal{O}(k^{\omega-1}m^{2})$ operations in $\mathbb{F}_{q^{d}}$. Solving the system of equations takes $\mathcal{O}(m)$ multiplications in $\mathbb{F}_{q^{d}}$. In total, the complexity is $\tilde{\mathcal{O}}(dk^{\omega-1}m^{2})$. ∎
@@ Appendix 0.A Normalizing matrices and characteristic polynomials
$\{U\in\mathbb{F}_{q}^{m\times m}\mid\operatorname{Tr}(U)=\operatorname{Tr}(U^{2})=0\}/\mathbb{F}_{q}^{\times}$ $\longrightarrow(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-3}\times\{0\}))/\mathbb{F}_{q}^{\times}$
@@
where the characteristic polynomial of $U$ is $t^{m}+a_{m-3}t^{m-3}+\dots+a_{1}t+a_{0}$ and $\mathbb{F}_{q}^{\times}$ acts on $\mathbb{F}_{q}^{m-2}$ via $\lambda\diamond(a_{m-3},\dots,a_{0})=(\lambda^{3}a_{m-3},\dots,\lambda^{m}a_{0})$.
@@
The set $(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-3}\times\{0\}))/\mathbb{F}_{q}^{\times}$ has
@@
The subset $\mathrm{Sep}_{q,m}\subset(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-3}\times\{0\}))/\mathbb{F}_{q}^{\times}$ of classes of separable polynomials has $\sim q^{m-3}$ elements.
@@
elements by the Serre bound [29, Théorème]. Since every element of $\mathbb{P}^{m-3}_{3,\dots,m}(\mathbb{F}_{q})$ has exactly $q-1$ preimages in $\mathbb{F}_{q}^{m-2}$ [23, Lem. 7], this means that there are $\mathcal{O}(mq^{m-4})$ classes of inseparable polynomials in $\mathbb{P}^{m-3}_{3,\dots,m}(\mathbb{F}_{q})$. Hence, by Lemma 14, $\mathrm{Sep}_{q,m}$ has $q^{m-3}-\mathcal{O}(mq^{m-4})\sim q^{m-3}$ elements.∎
@@
      $\mathbb{F}_{q}^{\times}/\mu_{d_{0}}(\mathbb{F}_{q})^{i_{1}}=\mathbb{F}_{q}^{\times}/\mu_{d_{0}/d_{1}}(\mathbb{F}_{q})$
@@
      If $d_{1}\neq 1$, continue with $d_{2}=\gcd(d_{1},i_{2})$. Stop after the $k$-th step if $k=\ell$ or $d_{k}=1$. Return $\chi^{\prime}=\lambda_{k}\diamond\chi$. The algorithm is summed up below.
@@
while $d\neq 1$ and $(a_{m-i-1},\dots,a_{1})\neq(0,\dots,0)$ do
@@
return $\lambda U$, $(\lambda^{3}a_{m_{3}},\dots,\lambda^{m}a_{0})$
@@
The complexity of Algorithm 7 is ${\mathcal{O}}(m\log(q)^{2})$. To reduce its impact on the complexity of our attack, one can precompute the normalization of the most frequent characteristic polynomials. For instance, on may compute in advance the suitable $\lambda$ for vectors $(a_{m-3},\dots,a_{0})$ such that $a_{m-3},a_{m-4}\neq 0$. Since 3 and 4 are coprime, a unique $\lambda$ is found knowing only $a_{m-3},a_{m-4}$: constructing a dictionary $\{(a_{m-3},a_{m-4}):\lambda\}$ allows to precompute the normalization for $(q-1)^{2}q^{m-4}\sim q^{m-2}$ elements of $\mathbb{F}_{q}^{m-2}$, that is, almost all of them.
@@ Appendix 0.B Proportion of codes with one-dimensional hull
Sendrier’s work [26] gives detailed results about the number of codes with a hull of given dimension. His results are proven in the case of codes inside $\mathbb{F}_{q}^{n}$ with the usual inner product. In our case however, we consider $\mathbb{F}_{q}^{m\times m}$ endowed with the bilinear form $(X,Y)\mapsto\operatorname{Tr}(XY)$. Denoting by $\sigma_{n,i}$ the number of totally isotropic $[n,k]_{q}$-codes for a given bilinear form, the number $A_{n,k,1}$-codes whose intersection with their orthogonal complement has dimension 1 is equal to [26, Theorem 2]
@@
where [${}{n}{k}$ ] is the Gaussian binomial coefficient which denotes the number of $k$-dimensional linear subspaces of $\mathbb{F}_{q}^{n}$. The proof of this result does not involve the nature of the considered non-degenerate bilinear form. Sendrier goes on to show [26, Theorem 3], using asymptotic results based on explicit values of $\sigma_{n,i}$ specific to a bilinear form of discriminant 1, that for $21\leqslant k\leqslant n/$ 2,
@@
There are different formulas of $\sigma_{n,i}$ given in [26, Theorem 1] depending on the remainders of $n,q$ modulo 2 and 4 and on the size of $k$. However, they are asymptotically equivalent, which yields this uniform result. For a bilinear form of different discriminant, these formulas are simply permuted; a general expression may be found in [8, IV, Proposition 3.5]. This does not change the asymptotic result above. Moreover, for $k\geqslant n/2$, the number of $[n,k]$-codes with one-dimensional hull is that of $[n,n-k]$-codes with one-dimensional hull, since the hull of a code $\mathcal{C}$ is exactly that of its dual. In particular, this means that for any $k$ such that $11\leqslant k\leqslant n-$ 1,